# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 3 Proofs of the Main Results
In this section, we prove our main theorems concerning the critical points (i.e., 1) and the error bound (i.e., 2) of Problem (2). To avoid interrupting the flow of the main proofs, we introduce key lemmas and propositions, with their detailed proofs deferred to Section 4. Based on these results, we present the complete proofs of 1 and 2.
3.1 Preliminary Results
To characterize the critical points and establish the error bound of Problem (2), we claim that it suffices to study the following problem:
$\min_{\bm{W}}\ G(\bm{W}):=\left\|\bm{W}_{L}\ldots\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right\|_{F}^{2}+\lambda\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2},$ (13)
where $\lambda:=\lambda_{1}\dots\lambda_{L}$. We compute the gradient of $G(\bm{W})$, which will be frequently used, as follows:
$\frac{1}{2}\nabla_{\bm{W}_{l}}G(\bm{W})=\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l-1:1}^{T}+\lambda\bm{W}_{l},\ \forall l\in[L].$ (14)
The critical point set of this problem is defined as
$\mathcal{W}_{G}:=\left\{\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L}):\nabla_{\bm{W}_{l}}G(\bm{W})=\bm{0},\ \forall l\in[L]\right\}.$ (15)
@@ 3.1 Preliminary Results
(i) $(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}$ if and only if $(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L})\in\mathcal{W}_{G}$.
(ii) Suppose that there exist constants $\epsilon,\kappa>0$ such that for all $\bm{Z}$ satisfying $\mathrm{dist}(\bm{Z},\mathcal{W}_{G})\leq\epsilon$, it holds that
@@
Then, for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W})\leq\epsilon/\sqrt{\lambda_{\max}}$, it holds that
$\mathrm{dist}(\bm{W},\mathcal{W})\leq\frac{\kappa\lambda}{\lambda_{\min}}\|\nabla F(\bm{W})\|_{F}.$
(iii) Let (5) be an SVD of $\bm{Y}$. It holds that $(\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{L})$ is a critical point of Problem (13) if and only if $(\bm{W}_{1}\bm{V}_{Y},\bm{W}_{2},\dots,\bm{U}_{Y}^{T}\bm{W}_{L})$ is a critical point of the following problem:
$\min_{\bm{W}}\ \left\|\bm{W}_{L}\ldots\bm{W}_{1}-\sqrt{\lambda}\bm{\Sigma}_{Y}\right\|_{F}^{2}+\lambda\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2}.$
@@
(i) Let $\bm{W}:=(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}$ be arbitrary. Using this and (15), we obtain
$\nabla_{\bm{W}_{l}}F(\bm{W})=2\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\bm{Y}\right)\bm{W}_{l-1:1}^{T}+2\lambda_{l}\bm{W}_{l}=\bm{0},\ \forall l\in[L].$ (17)
Now, let $\hat{\bm{W}}:=(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L})$. For each $l\in[L]$, we compute
@@
Therefore, we have $\hat{\bm{W}}\in\mathcal{W}_{G}$. Conversely, let $\hat{\bm{W}}\in\mathcal{W}_{G}$ be arbitrary. Using the same argument, we have $\bm{W}\in\mathcal{W}$.
(ii) Using (i), we express $\mathcal{W}_{G}$ as
@@
$\mathcal{W}_{G}=\left\{(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L}):(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}\right\}.$
Now, let $\bm{W}:=\left(\bm{W}_{1},\dots,\bm{W}_{L}\right)$ be arbitrary and $\bm{Z}:=(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L})$. Then, let $(\sqrt{\lambda_{1}}\hat{\bm{W}}_{1}^{*},\dots,\sqrt{\lambda_{L}}\hat{\bm{W}}_{L}^{*})$ with $\hat{\bm{W}}^{*}\in\mathcal{W}$ be such that $\mathrm{dist}^{2}(\bm{Z},\mathcal{W}_{G})=\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}-\hat{\bm{W}}^{*}_{l}\|_{F}^{2}$ and $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}$ be such that $\mathrm{dist}^{2}(\bm{W},\mathcal{W})=\sum_{l=1}^{L}\|\bm{W}_{l}-\bm{W}_{l}^{*}\|_{F}^{2}$. We obtain

### v3

@@ 3 Proofs of the Main Results
In this section, we present the proofs of our main results on the critical point set (i.e., 1) in Section 3.1 and the error bound (i.e., 2) for Problem (1) in Section 3.2. To facilitate readability, we summarize the notations used in this section in Table 1.
3.1 Proofs of the Critical Point Set
To characterize the critical points and establish the error bound for Problem (1), we claim that it suffices to study the following problem:
$\min_{\bm{W}}\ G(\bm{W}):=\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right\|_{F}^{2}+\lambda\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2},$ (12)
where $\lambda:=\prod_{l=1}^{L}\lambda_{l}$, as defined in (7) and (8). We compute the gradient of $G(\bm{W})$ as follows:
$\nabla_{l}G(\bm{W})=2\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l-1:1}^{T}+2\lambda\bm{W}_{l},\ \forall l\in[L].$ (13)
The critical point set of Problem (12) is defined as
$\mathcal{W}_{G}:=\left\{\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L}):\nabla G(\bm{W})=\bm{0}\right\}.$
Notation Eq. Notation Eq. Notation Eq. Notation Eq. Notation Eq. Notation Eq.
$p$ (45) $p_{Y},r_{Y}$ (4) $s_{i},h_{i}$ (4) $t_{i},g_{i}$ (45) $\varphi(x)$ (85) $G(\bm{W})$ (12)
$\lambda$ (7,8) $\delta_{y}$ (18) $\delta_{\sigma},\mathcal{\bm{Y}}$ (28) $\mathcal{\bm{A}}$ (26) $\mathcal{\bm{B}}$ (27) $\mathcal{\bm{A}}_{\text{sort}}$ (29)
$\eta_{1}$ (56) $\eta_{2}$ (84) $\eta_{3},\eta_{4}$ (88) $\eta_{5}$ (90) $L_{G}$ (144) $c_{1}$ (42)
$c_{2}$ (57) $c_{3}$ (74) $c_{4}$ (90) $c_{5}$ (95) $\delta_{1},\delta_{2}$ (87) $\sigma^{*}_{\min},\sigma_{\max}^{*},r_{\sigma}$ (34)
$\epsilon_{\bm{0}},\kappa_{\bm{0}}$ (31) $\epsilon_{\bm{\sigma}^{*}},\kappa_{\bm{\sigma}^{*}}$ (114) $\mathcal{W}_{\bm{\sigma}^{*}}$ (19,29) $\kappa,\epsilon$ (125) $\kappa_{1},\epsilon_{1}$ (126) $d_{\max}$ (18)
Table 1: The explicit expressions of the notations in the paper
@@ 3.1 Proofs of the Critical Point Set
(i) $(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}$ if and only if $(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L})\in\mathcal{W}_{G}$.
(ii) Let $\bm{Y}=\bm{U}_{Y}\bm{\Sigma}_{Y}\bm{V}_{Y}^{T}$ be an SVD of $\bm{Y}$. Then $(\bm{W}_{1},\ldots,\bm{W}_{L})$ is a critical point of Problem (12) if and only if $(\bm{W}_{1}\bm{V}_{Y},\,\bm{W}_{2},\,\ldots,\,\bm{W}_{L-1},\,\bm{U}_{Y}^{T}\bm{W}_{L})$ is a critical point of $\min_{\bm{W}}\,\big\|\,\bm{W}_{L:1}-\sqrt{\lambda}\bm{\Sigma}_{Y}\big\|_{F}^{2}+\lambda\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2}.$
(iii) Suppose that there exist constants $\epsilon,\kappa>0$ such that for all $\bm{Z}$ satisfying $\mathrm{dist}(\bm{Z},\mathcal{W}_{G})\leq\epsilon$, we have
@@
Then for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon/\sqrt{\lambda_{\max}}$, it holds that
$\mathrm{dist}\left(\bm{W},\mathcal{W}_{F}\right)\leq({\kappa\lambda}/{\lambda_{\min}})\|\nabla F(\bm{W})\|_{F}.$
@@
(i) and (ii) have been proved in [6, Lemma 3.1]. Now, we prove (iii). Using (i), we have
$\mathcal{W}_{G}=\left\{(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L}):(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}\right\}.$
Let $\bm{W}=\left(\bm{W}_{1},\dots,\bm{W}_{L}\right)$ be arbitrary and $\hat{\bm{W}}=(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L})$. Let $(\sqrt{\lambda_{1}}\hat{\bm{W}}_{1}^{*},\dots,\sqrt{\lambda_{L}}\hat{\bm{W}}_{L}^{*})$ with $(\hat{\bm{W}}_{1}^{*},\dots,\hat{\bm{W}}_{L}^{*})\in\mathcal{W}_{F}$ be such that $\mathrm{dist}^{2}(\hat{\bm{W}},\mathcal{W}_{G})=\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}-\hat{\bm{W}}^{*}_{l}\|_{F}^{2}$, and let $(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{F}$ be such that $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})=\sum_{l=1}^{L}\|\bm{W}_{l}-\bm{W}_{l}^{*}\|_{F}^{2}$. We have
$\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})$ $=\sum_{l=1}^{L}\|\bm{W}_{l}-\bm{W}_{l}^{*}\|_{F}^{2}\geq\frac{\sum_{l=1}^{L}\|\sqrt{\lambda_{l}}\bm{W}_{l}-\sqrt{\lambda_{l}}\bm{W}_{l}^{*}\|_{F}^{2}}{\lambda_{\max}}$
$\geq\frac{1}{\lambda_{\max}}\mathrm{dist}^{2}(\hat{\bm{W}},\mathcal{W}_{G}).$ (15)
Using this and $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon/\sqrt{\lambda_{\max}}$ yields $\mathrm{dist}(\hat{\bm{W}},\mathcal{W}_{G})\leq\epsilon$, and thus (14) holds. Next, we have
$\textstyle\mathrm{dist}^{2}(\hat{\bm{W}},\mathcal{W}_{G})\!=\!\!\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}-\hat{\bm{W}}^{*}_{l}\|_{F}^{2}\geq\!\lambda_{\min}\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F}),$ (16)
@@
$\qquad\qquad\qquad\qquad\ =\frac{\lambda}{\sqrt{\lambda_{l}}}\nabla_{l}F(\bm{W}),\ \forall l\in[L].$ (17)
Therefore, we have
@@
Using (i) and (iii) of the above lemma, it suffices to establish an error bound for Problem (12) in the rest of this section. By Lemma 1(ii), we assume without loss of generality that $\bm{Y}=\bm{\Sigma}_{Y}$ from now on, where $\bm{\Sigma}_{Y}$ is defined in (3) and its diagonal entries are partitioned as in (4). Moreover, we define
$\delta_{y}:=\min\left\{|y_{s_{i}}-y_{s_{i+1}}|:i\in[p_{Y}]\right\},\quad d_{\max}:=\max\{d_{0},d_{1},\ldots,d_{L}\}.$ (18)
Using [6, Proposition 3.3] and Assumption 1, it is straightfoward to verify that the critical point set of Problem (12) admits the following closed-form expression:
Fact 1.
Suppose that Assumption 1 holds. The critical point set of Problem (12) is
$\mathcal{W}_{G}=\bigcup_{(\bm{\sigma},\bm{\Pi})\in\mathcal{B}}\mathcal{W}_{(\bm{\sigma},\bm{\Pi})},$ (19)
where
$\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}:=\left\{\bm{W}:\begin{array}[]{l}\bm{\Sigma}_{l}=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\ \forall l\in[L],\\ \bm{W}_{1}=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{\Pi},\bm{I}\right)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ \bm{W}_{l}=\bm{Q}_{l+1}\bm{\Sigma}_{l}\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\ \bm{Q}_{l}\in\mathcal{O}^{d_{l-1}},\ l=2,\dots,L,\\ \bm{W}_{L}=\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\mathrm{BlkD}\left(\bm{\Pi}^{T},\bm{I}\right)\bm{\Sigma}_{L}\bm{Q}_{L}^{T},\\ \bm{O}_{i}\in\mathcal{O}^{h_{i}},\ \forall i\in[p_{Y}],\ \bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\ \widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}\end{array}\right\},$ (25)
$\mathcal{A}:=\left\{\bm{a}\in\mathbb{R}^{d_{\min}}:\ a_{i}^{2L-1}-\sqrt{\lambda}y_{i}a_{i}^{L-1}+\lambda a_{i}=0,\ a_{i}\geq 0,\ \forall i\in[d_{\min}]\right\},$ (26)
$\mathcal{B}:=\left\{(\bm{\sigma},\bm{\Pi})\in\mathbb{R}^{d_{\min}}\times\mathcal{P}^{d_{\min}}:\bm{a}\in\mathcal{A},\ \bm{\sigma}=\bm{\Pi}\bm{a},\ \sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{d_{\min}}\right\}.$ (27)
Note that the permutation matrix $\bm{\Pi}$ can be absorbed into $\bm{\Sigma}_{l}$ and $\bm{Q}_{l}$ to simplify the above expression. With this observation, and by invoking Lemma 1, we are now ready to prove 1 as follows:
Proof.
By (ii) of Lemma 1, without loss of generality, we may assume that $\bm{Y}=\bm{\Sigma}_{Y}$, i.e., $\bm{U}_{Y}=\bm{I}$ and $\bm{V}_{Y}=\bm{I}$. Let $(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}$ be arbitrary. It follows from 1 and Lemma 1 that there exists $(\bm{\sigma},\bm{\Pi})\in\cal$ B such that $(\bm{W}_{1},\dots,\bm{W}_{L})$ takes a form of (25) with $\bm{\Sigma}_{l}=\mathrm{BlkDiag}(\mathrm{diag}(\bm{\sigma})/\sqrt{\lambda_{l}},\bm{0})$. According to (26) and (27), there exists $\bm{a}\in\cal A$ such that $\bm{\Pi}\mathrm{diag}(\bm{a})\bm{\Pi}^{T}=\mathrm{diag}(\bm{\sigma})$. Then, let
$\bm{\Sigma}_{l}^{\prime}:=\mathrm{BlkDiag}(\bm{\Pi}^{T},\bm{I}_{d_{l}-d_{\min}})\bm{\Sigma}_{l}\mathrm{BlkDiag}(\bm{\Pi},\bm{I}_{d_{l-1}-d_{\min}})=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{a})/\sqrt{\lambda_{l}},\bm{0}\right),\ \forall l\in[L],$
Therefore, we obtain that $(\bm{W}_{1},\dots,\bm{W}_{L})$ can be rewriten in the form of (5) with $\bm{\Sigma}_{l}=\bm{\Sigma}_{l}^{\prime}$, $\bm{Q}_{l}=\bm{Q}_{l}^{\prime}$. Conversely, let $(\bm{W}_{1},\dots,\bm{W}_{L})$ be of the form of (5) for some $\bm{\sigma}=(\sigma_{1},\ldots,\sigma_{r_{Y}},0,\ldots 0)\in\mathbb{R}^{d_{\min}}$. Then, we choose a permutation matrix $\bm{\Pi}\in\mathcal{P}^{d_{\min}}$ such that $\bm{\sigma}^{\prime}=\bm{\Pi}\bm{\sigma}\ \text{satisfies}\ \sigma^{\prime}_{1}\geq\sigma^{\prime}_{2}\geq\dots\geq\sigma^{\prime}_{d_{\min}}.$ Let
$\bm{\Sigma}^{\prime}_{l}:=\mathrm{BlkDiag}(\bm{\Pi},\bm{I}_{d_{l}-d_{\min}})\bm{\Sigma}_{l}\mathrm{BlkDiag}(\bm{\Pi}^{T},\bm{I}_{d_{l-1}-d_{\min}})=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\sigma}^{\prime})/\sqrt{\lambda_{l}},\bm{0}\right),\ \forall l\in[L],$
Then, we obtain that $(\bm{W}_{1},\dots,\bm{W}_{L})$ can be rewriten in the form of (25) with $\bm{\Sigma}_{l}=\bm{\Sigma}_{l}^{\prime}$, $\bm{Q}_{l}=\bm{Q}_{l}^{\prime}$. This, together with 1 and item (i) and (ii) of Lemma 1, implies $(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}$. ∎
3.2 Analysis of the Error Bound
In this subsection, we establish the error bound for Problem (12) based on 1. In Section 3.2.1, we first present preliminary results for our analysis. In Section 3.2.2, we provide the main proofs of the error bound. For ease of exposition, we define
@@ 3.2 Analysis of the Error Bound
$\mathcal{Y}:=\bigcup_{i\in[d_{\min}]}\left\{\sigma\geq 0:\sigma^{2L-1}+\lambda\sigma-\sqrt{\lambda}y_{i}\sigma^{L-1}=0\right\},\ \delta_{\sigma}:=\min\left\{|x-y|:x\neq y\in\mathcal{Y}\right\}.$ (28)
The following result elucidates the structure of the collection $\{\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}\}_{(\bm{\sigma},\bm{\Pi})\in\mathcal{B}}$.
Proposition 1.
Let $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$ and $(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})\in\mathcal{B}$ be arbitrary. The following statements hold:
(i) It holds that $\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}=\mathcal{W}_{(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})}$ if and only if $\bm{\sigma}=\bm{\sigma}^{\prime}$.
(ii) If $\bm{\sigma}\neq\bm{\sigma}^{\prime}$, it holds that $\mathrm{dist}\left(\mathcal{W}_{(\bm{\sigma},\bm{\Pi})},\mathcal{W}_{(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})}\right)\geq\delta_{\sigma}.$
We defer the proof of this proposition to Appendix B.1. This proposition demonstrates that for any pair $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$ and $(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})\in\mathcal{B}$, if $\bm{\sigma}\neq\bm{\sigma}^{\prime}$, $\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}$ is well separated from $\mathcal{W}_{(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})}$, and otherwise they are identical. Therefore, with a slight abuse of notation, we write $\mathcal{W}_{\bm{\sigma}}:=\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}$ for any $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$. Thus, we can express the set of critical points of Problem (12) as follows:
3.2.1 Preliminary Results
According to 1 and (29), for any $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{G})\leq\delta_{\sigma}/2$, there exists a $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}$ such that $\mathrm{dist}(\bm{W},\mathcal{W}_{G})=\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}).$ This observation simplifies our analysis, as it suffices to bound $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ for each $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}$. Note that $\bm{0}\in\mathcal{A}_{\rm sort}$. For this case, we directly show the error bound as follows:
Proposition 2.
Suppose that Assumptions 1 and 2 hold. Then for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{0}})\leq\epsilon_{\bm{0}}$, we have
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{0}})\leq\kappa_{\bm{0}}\,\bigl\|\nabla G(\bm{W})\bigr\|_{F},$ (30)
where
$(\epsilon_{\bm{0}},\,\kappa_{\bm{0}}):=\begin{cases}\Biggl(\Bigl(\frac{\sqrt{\lambda}}{2(\sqrt{\lambda}+y_{1})}\min\!\Bigl\{\min\limits_{i\in[s_{p_{Y}}]}\bigl|\lambda-y_{i}^{2}\bigr|,\,\lambda\Bigr\}\Bigr)^{\frac{1}{2}},\;\frac{2(\sqrt{\lambda}+y_{1})}{\sqrt{\lambda}\min\!\Bigl\{\min\limits_{i\in[s_{p_{Y}}]}\bigl|\lambda-y_{i}^{2}\bigr|,\,\lambda\Bigr\}}\Biggr),\!\!&\!\!\!\!L=2,\\ \Biggl(\min\!\left\{\left(\frac{\lambda}{3}\right)^{{1}/{(2L-2)}},\;\left(\frac{\sqrt{\lambda}}{3y_{1}}\right)^{{1}/{(L-2)}}\right\},\;\frac{3\sqrt{L}}{2\lambda}\Biggr),\!\!&\!\!\!\!L\geq 3.\end{cases}$ (31)
Proof.
Let’s begin by focusing on $L=2$. Using 1, we have $\mathcal{W}_{\bm{0}}=\{(\bm{0},\bm{0})\}$ and $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})=\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}.$ Without loss of generality, assume that $\|\bm{W}_{1}\|_{F}\geq\|\bm{W}_{2}\|_{F}$. We have
$\nabla_{1}G(\bm{W})=2\bm{W}_{2}^{T}\left(\bm{W}_{2}\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right)+2\lambda\bm{W}_{1},$
$\nabla_{2}G(\bm{W})=2\left(\bm{W}_{2}\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{1}^{T}+2\lambda\bm{W}_{2}.$ (32)
Using the first equation in (32), we have
$\frac{\sqrt{\lambda}}{2}\|\nabla_{1}G(\bm{W})\|_{F}$ $\geq\left\|\lambda\bm{W}_{2}^{T}\bm{Y}-\lambda^{\frac{3}{2}}\bm{W}_{1}\right\|_{F}-\sqrt{\lambda}\|\bm{W}_{2}^{T}\bm{W}_{2}\bm{W}_{1}\|_{F}$
Multiplying $\bm{Y}^{T}$ on the both sides of the second equation in (32), together with the triangle inequality, yields
$\frac{1}{2}\left\|\bm{Y}^{T}\nabla_{2}G(\bm{W})\right\|_{F}\geq\|\sqrt{\lambda}\bm{Y}^{T}\bm{Y}\bm{W}_{1}^{T}-\lambda\bm{Y}^{T}\bm{W}_{2}\|_{F}-\|\bm{Y}^{T}\bm{W}_{2}\bm{W}_{1}\bm{W}_{1}^{T}\|_{F},$
which implies
$\frac{y_{1}}{2}\|\nabla_{2}G(\bm{W})\|_{F}\geq\|\sqrt{\lambda}\bm{Y}^{T}\bm{Y}\bm{W}_{1}^{T}-\lambda\bm{Y}^{T}\bm{W}_{2}\|_{F}-y_{1}\|\bm{W}_{2}\|_{F}\|\bm{W}_{1}\|_{F}^{2}.$
Summing up this inequality with (3.2.1) and using the triangle inequality, we obtain
$\frac{\sqrt{\lambda}+y_{1}}{2}\|\nabla G(\bm{W})\|_{F}$ $\geq\frac{\sqrt{\lambda}}{2}\|\nabla_{1}G(\bm{W})\|_{F}+\frac{y_{1}}{2}\left\|\nabla_{2}G(\bm{W})\right\|_{F}$
$\geq\sqrt{\lambda}\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\lambda-y_{i}^{2}\right|,\lambda\right\}\|\bm{W}_{1}\|_{F}-\left(\sqrt{\lambda}+y_{1}\right)\|\bm{W}_{1}\|_{F}^{3}$